# Efficient Quantum Pseudorandomness from Hamiltonian Phase States

### v1

@@ Abstract
Quantum pseudorandomness has found applications in many areas of quantum information, ranging from entanglement theory, to models of scrambling phenomena in chaotic quantum systems, and, more recently, in the foundations of quantum cryptography. Kretschmer (TQC ’21) showed that both pseudorandom states and pseudorandom unitaries exist even in a world without classical one-way functions. To this day, however, all known constructions require classical cryptographic building blocks which are themselves synonymous with the existence of one-way functions, and which are also challenging to realize on realistic quantum hardware.
@@ Acknowledgements
The authors would like to thank Aram Harrow, David Gross, Eli Goldin, Fermi Ma, Henry Yuen, Kabir Tomer, Luowen Qian, Natalie Parham, Ran Canetti, Saachi Mutreja, Soonwon Choi, Thiago Bergamaschi, Uma Girish and Vinod Vaikuntanathan for useful discussions.
@@ 3.4 The Haar Measure and $t$-designs
$\left|\left|\underset{\psi\sim\nu}{\mathbb{E}}(|\psi\rangle\langle\psi|)^{\otimes t}-\underset{\ket{\psi}\sim\mathsf{Haar}(d)}{\mathbb{E}}(|\psi\rangle\langle\psi|)^{\otimes t}\right|\right|_{1}\leq\varepsilon.$ (3.6)
@@ 5.2 Hamiltonian Phase States Form Approximate State Designs
We provide a proof of Theorem 5.16 at the end of this subsection. The proof of Theorem 5.16 is remarkably simple in comparison to the derivations of similar results for random quantum circuits [chen2024incompressibility, brandao2016local, haferkamp2022random]. Additionally, the constants in Theorem 5.16 are unusually small: In stark contrast the constants in these results are north of $10^{13}$. A similar result was obtained in Ref. [haah2024efficient] for the related random Pauli rotations $e^{\mathrm{i}\theta P}$ for a random $\theta\in(0,2\pi]$ and a random Pauli string $P$.
@@ 5.4 Discussion of Full Quantumness
One way that researchers get around this is to take some “idealized” version of the assumption, and show that relative to some oracle, the idealized version exists while one-way functions do not. The oracle separations shown in [kretschmer2021quantum, kretschmer2023quantum] already prove this for us. Namely, if we assume that our circuits actually output Haar random states, then with an additional $\mathsf{PSPACE}$ oracle, $\mathsf{BQP}=\mathsf{QMA}$, but our idealized Haar random states are genuinely Haar random. However, using this as evidence that our assumption is fully-quantum is barely better than simply stating it to be so, since our argument essentially reduced to “nothing that claims to output Haar random-looking states can imply one-way functions”. However, the fact that our ensembles are actually $t$-designs for a very high $t$ should make one-way very hard to build from these states in practice. Indeed, examining the proof of [kretschmer2021quantum], we can see that if we instead take a family of $t$-designs instead of Haar random unitaries, any $\mathsf{BQP}$ algorithm that calls the oracle fewer than $t$ times will have a similar concentration as the Haar measure, and thus will be simulatable by the $\mathsf{PSPACE}$ oracle (without access to the $t$-design). This implies that any construction of one-way functions that directly uses our assumption must make many (more than $t=\sqrt{m/n}$) calls to our oracle, implying that any such construction cannot be simple.
@@ 6.5 Quantum Pseudoentanglement
    |$\ket{\Phi_{\boldsymbol{\omega}}^{\mathbf{B}}}=\exp\left(\mathrm{i}\,\sum_{i=1}^{k}\theta_{i}\bigotimes_{j=1}^{n}\mathsf{Z}^{\mathbf{B}_{ij}}\right)H^{\otimes n}\ket{0^{n}}.$
@@ 6.6 Pseudorandom Unitaries
  • (Pseudorandomness) The unitary $U_{k}$, for a randomly chosen key $k\sim\mathcal{K}$, is computationally indistinguishable from a Haar random unitary $U\sim\mathsf{Haar}(2^{n})$. In other words, for any efficient quantum algorithm $\mathcal{A}$, it holds that
    $\vline\,\underset{k\sim\mathcal{K}}{\Pr}[\mathcal{A}^{U_{k}}(1^{n})=1]-\underset{U\sim\mathsf{Haar}(2^{n})}{\Pr}[\mathcal{A}^{U}(1^{n})=1]\,\vline\,\leq\,\mathsf{negl}(n)\,.$
@@ 6.6.1 Security of the $\mathsf{FHFHFC}$ pseudo-random unitary
Let $\mathsf{F}$ be an “ideal HPS” phase oracle, which acts as follows for a tuple of angles $\Theta=(\theta_{z})_{z}$:
@@
Here, each angle $\theta_{z}$ will be randomly drawn from a discrete set $\{2\pi k/q\}_{k=0}^{q-1}$. In the following, we will choose $q=2^{n}$. We note that any exponentially large $q$ suffices. Note that we call this an ideal HPS because it involves sampling an exponential number of real phases, instead of only a polynomial number of them. We claim that the following unitary is pseudo-random, for two ideal HPS instances $\mathsf{F}_{1}$ and $\mathsf{F}_{2}$ with independently sampled angles:
@@
Before we begin with the proof, we comment on the construction itself: Why take three rounds of $\mathsf{F}$? Intuitively, applying a single random function $\mathsf{F}$ records the input to the function, as noted in [zhandry2019record]. Therefore, we can see that each of the three random phase oracles has a distinct “job”: The first one records the input, and the last one records the output. The middle one records the relationship between the input and output, as we will see later. In this manner, the $\mathsf{FHFHF}$ unitary implements the path recording oracle from [ma2024pseudorandom]. We also note that previous work [nakata2017unitary] showed a similar transition at $3$-rounds of random diagonal and Hadamard unitaries, which we will comment on later.
@@
[…]
@@
$\frac{1}{2^{3n/2}}\sum_{\Theta_{1},\Theta_{2},\Theta_{3}}\ket{\Theta_{1},\Theta_{2},\Theta_{3}}\,.$
@@
Since the view of the adversary does not contain the purifying register, when we trace them out we will get a uniform mixture of states corresponding to $\mathsf{FHF}$ for uniformly random angles $\Theta_{1}$, $\Theta_{2}$, and $\Theta_{3}$. This corresponds exactly to choosing $\Theta_{1}$, $\Theta_{2}$, and $\Theta_{3}$ uniformly at random. ∎
In the remainder of this section, we will write $\mathop{\mathbb{E}}_{\Theta_{1},\Theta_{2},\Theta_{3}}\ket{\Theta_{1},\Theta_{2},\Theta_{3}}$, but this should be taken to be the normalized vector, as written in Lemma 6.18. We now define a family of orthogonal states that correspond to a database $R$. We some new notation here to make the remaining equations less cumbersome. For a set $S$ of $n$-bit strings and $n$-bit string $z$, define $S_{z}$ to be:
@@
$\mathrm{Im}(R)$ =$=\{y:(x,y)\in R\}$
$\mathrm{Zip}(R)$ =$=\{x\oplus y:(x,y)\in R\}\,.$
Definition 6.19 (FHFHF-relation states).
|$\ket{\mathsf{fhfhf}_{R}}=\mathbb{E}_{\Theta^{(1)},\Theta^{(2)},\Theta^{(3)}}\exp\left(i\sum_{z}\theta^{(1)}_{z}\mathrm{Dom}(R)_{z}+\theta^{(2)}_{z}\mathrm{Im}(R)_{z}+\theta_{z}^{(3)}\mathrm{Zip}(R)_{z}\right)\ket{\Theta_{1}}\otimes\ket{\Theta_{2}}\otimes\ket{\Theta_{3}}\,.$
We show that these states form an orthogonal basis. In order to prove this, we first note the following lemma about the Welsh-Hadamard encoding.
@@
Let $R$ and $R^{\prime}$ be two different bijective relations of size at most $2^{n-1}$ elements. Then |$\ket{\mathsf{fhfhf}_{R}}$ and |$\ket{\mathsf{fhfhf}_{R^{\prime}}}$ are orthogonal.
@@
First, consider two relations such that their domains differ, and let $z^{*}$ be the index implied by the previous lemma for the sets $\mathrm{Dom}(R)$ and $\mathrm{Dom}(R^{\prime})$. We proceed by evaluating the inner product of the two relation states.
[…]
$\hskip 68.2866pt=0\,.$
Here, we use the following normalized integral over a continuous angle $\theta\in(0,2\pi]$:
$\mathbb{E}_{\theta}e^{i\theta x}=\begin{cases}1\quad x=0\\ 0\quad\text{otherwise}\end{cases}$
Moreover, it is easy to check tht $\mathbb{E}_{k}e^{\mathrm{i}2\pi k/q}=\mathbb{E}_{\theta}e^{i\theta x}$ for an integer $x$ with $|x|<q$. Clearly,
$|\mathrm{Dom}(R)_{z}-\mathrm{Dom}(R^{\prime})_{z}|\leq 2t<q~{}$
and similarly for $\mathrm{Im}(R)_{z}-\mathrm{Im}(R^{\prime})_{z}$ and $\mathrm{Zip}(R)_{z}-\mathrm{Zip}(R^{\prime})_{z}$. Moreover, by Lemma 6.20, the term multiplying $\theta_{z^{*}}$ is non-zero, which means that expectation in the product will vanish.
By a similar argument, if the two relations differ on their image, then there will be another $z^{*}$ we can take. Finally, if the two relations have the same domain and image, they must differ on the XOR of their values, so there will again be a $z^{*}$ that can be taken to make the inner product 00. ∎
@@
$\mathsf{FHFHFO}\ket{x}\ket{\mathsf{fhf}_{R}}=\frac{1}{\sqrt{2^{n}-|R|}}\sum_{y}\ket{y}\ket{\mathsf{fhfhf}_{R\cup\{(x,y)\}}}\,.$
@@
|$\ket{x}\mathbb{E}_{\Theta_{1},\Theta_{2},\Theta_{3}}\alpha_{R}^{\Theta_{1},\Theta_{2},\Theta_{3}}\exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}\right)\ket{\Theta_{1},\Theta_{2},\Theta_{3}}\,.$
@@
$\sum_{y\in\{0,1\}^{n}}(-1)^{\braket{x,y}}\ket{y}\mathbb{E}_{\Theta_{1},\Theta_{2},\Theta_{3}}\alpha_{R}^{\Theta_{1},\Theta_{2}}\exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}\right)\ket{\Theta_{1},\Theta_{2}}\,.$
@@
$\sum_{y\in\{0,1\}^{n}}(-1)^{\braket{x,y}}\ket{y}\mathbb{E}_{\Theta_{1},\Theta_{2},\Theta_{3}}\alpha_{R}^{\Theta_{1},\Theta_{2},\Theta_{3}}\exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}+\theta_{z}^{(2)}(-1)^{\braket{y,z}}\right)\ket{\Theta_{1},\Theta_{2},\Theta_{3}}\,.$

### v2

@@ Abstract
Quantum pseudorandomness has found applications in many areas of quantum information, ranging from entanglement theory, to models of scrambling phenomena in chaotic quantum systems, and, more recently, in the foundations of quantum cryptography. Kretschmer (TQC ’21) showed that both pseudorandom states and pseudorandom unitaries exist even in a world without classical one-way functions. To this day, however, all known constructions require classical cryptographic building blocks which are themselves synonymous with the existence of one-way functions, and which are also challenging to implement on realistic quantum hardware.
@@ Acknowledgements
The authors would like to thank Fermi Ma for pointing out a bug in an earlier version of the paper related to the pseudorandom states construction, and for helpful discussions. The authors would also like to thank Aram Harrow, David Gross, Eli Goldin, Henry Yuen, Kabir Tomer, Luowen Qian, Natalie Parham, Ran Canetti, Saachi Mutreja, Soonwon Choi, Thiago Bergamaschi, Uma Girish and Vinod Vaikuntanathan for useful discussions.
@@ 3.4 The Haar Measure and $t$-designs
$\left|\left|\underset{\psi\sim\nu}{\mathbb{E}}(|\psi\rangle\langle\psi|)^{\otimes t}-\underset{\ket{\psi}\sim\mu(d)}{\mathbb{E}}(|\psi\rangle\langle\psi|)^{\otimes t}\right|\right|_{1}\leq\varepsilon.$ (3.6)
@@
Here $\mu_{D}$ is the Haar measure (unique left- and right- invariant measure) over diagonal unitary matrices.
@@ 5.2 Hamiltonian Phase States Form Approximate State Designs
We provide a proof of Theorem 5.16 in Appendix A. The proof of Theorem 5.16 is remarkably simple in comparison to the derivations of similar results for random quantum circuits [chen2024incompressibility, brandao2016local, haferkamp2022random]. Additionally, the constants in Theorem 5.16 are unusually small: In stark contrast the constants in these results are north of $10^{13}$. A similar result was obtained in Ref. [haah2024efficient] for the related random Pauli rotations $e^{\mathrm{i}\theta P}$ for a random $\theta\in(0,2\pi]$ and a random Pauli string $P$.
@@ 5.4 Discussion of Full Quantumness
One way that researchers get around this is to take some “idealized” version of the assumption, and show that relative to some oracle, the idealized version exists while one-way functions do not. The oracle separations shown in [kretschmer2021quantum, kretschmer2023quantum] already prove this for us. Namely, if we assume that our circuits actually output Haar random states, then with an additional $\mathsf{PSPACE}$ oracle, $\mathsf{BQP}=\mathsf{QMA}$, but our idealized Haar random states are genuinely Haar random. However, using this as evidence that our assumption is fully-quantum is barely better than simply stating it to be so, since our argument essentially reduced to “nothing that claims to output Haar random-looking states can imply one-way functions”. However, the fact that our ensembles are actually $t$-designs for a very high $t$ should make one-way functions very hard to build from these states in practice. Indeed, examining the proof of [kretschmer2021quantum], we can see that if we instead take a family of $t$-designs instead of Haar random unitaries, any $\mathsf{BQP}$ algorithm that calls the oracle fewer than $t$ times will have a similar concentration as the Haar measure, and thus will be simulatable by the $\mathsf{PSPACE}$ oracle (without access to the $t$-design). This implies that any construction of one-way functions that directly uses our assumption must make many (more than $t=\sqrt{m/n}$) calls to our oracle, implying that any such construction cannot be simple.
@@ 6.5 Quantum Pseudoentanglement
    |$\ket{\Phi_{\boldsymbol{\omega}}^{\mathbf{B}}}=\exp\left(\mathrm{i}\,\sum_{i=1}^{k}\omega_{i}\bigotimes_{j=1}^{n}\mathsf{Z}^{\mathbf{B}_{ij}}\right)H^{\otimes n}\ket{0^{n}}.$
@@ 6.6 Pseudorandom Unitaries
  • (Pseudorandomness) The unitary $U_{k}$, for a randomly chosen key $k\sim\mathcal{K}$, is computationally indistinguishable from a Haar random unitary $U\sim\mu(2^{n})$. In other words, for any efficient quantum algorithm $\mathcal{A}$, it holds that
    $\vline\,\underset{k\sim\mathcal{K}}{\Pr}[\mathcal{A}^{U_{k}}(1^{n})=1]-\underset{U\sim\mu(2^{n})}{\Pr}[\mathcal{A}^{U}(1^{n})=1]\,\vline\,\leq\,\mathsf{negl}(n)\,.$
@@ 6.6.1 Security of the $\mathsf{FHFHFC}$ pseudo-random unitary
Let $\mathsf{F}$ a uniformly random diagonal matrix. We can think of $\mathsf{F}$ being an “ideal HPS” phase oracle, which acts as follows for a tuple of angles $\theta=(\theta_{z})_{z}$:
@@
Here, each angle $\theta_{z}$ will be randomly drawn from a discrete set $\{2\pi k/q\}_{k=0}^{q-1}$. In the following, we will choose $q=2^{n}$. We note that any exponentially large $q$ suffices. We make two observations before continuing with the proof. First, we call this an ideal HPS because it involves sampling an exponential number of real phases, instead of only a polynomial number of them. Second, we note that these matrices are identically distributed to uniformly random diagonal matrices with complex phases, but we phrase it as an ideal HPS instance to motivate our conjectured pseudo-random unitary construction from HPS.
Remark 6.16.
As a matter of notation, we will write $\theta$, or $\theta^{(1)}$, to mean the tuple of angles $(\theta_{1},\ldots,\theta_{2^{n}})$ (or with the appropriate superscript). The association between the list of angles associated with a $\theta$ will be clear from context.
We claim that the following unitary is pseudo-random, for three uniformly random diagonal unitaries with complex phases, $\mathsf{F}_{1}$, $\mathsf{F}_{2}$ and $\mathsf{F}_{3}$:
@@
We call this the $\mathsf{FHFHFC}$-ensemble, and by the previous observation, this is identical to the proposed construction from [ji2018pseudorandom]. The following is the main theorem of this section.
Theorem 6.17.
Let $n\in\mathbb{N}$ be the number of qubits as input to a quantum adversary. Then for all $t$-query adversaries $\mathcal{A}_{t}^{(\cdot)}$ making queries to a $n$ qubit oracle, the following holds
$\mathrm{td}\left(\mathbb{E}_{U\sim\mu(2^{n})}\left[\mathcal{A}_{t}^{U}(\ket{0})\right],\mathbb{E}_{\mathcal{O}\sim\mathsf{FHFHFC}}\left[\mathcal{A}_{t}^{\mathcal{O}}(\ket{0})\right]\right)\leq O\left(\sqrt{\frac{t^{3}}{2^{n}}}\right)\,.$
@@
Before we begin with the proof, we comment on the construction itself: Why take three rounds of $\mathsf{F}$? Intuitively, applying a single random function $\mathsf{F}$ records the input to the function, as noted in [zhandry2019record]. Therefore, we can see that each of the three random phase oracles has a distinct “job”: The first one records the input, and the last one records the output. The middle one records the relationship between the input and output, as we will see later. Where our proof differs from the proof of [ma2024pseudorandom] is that the middle unitary is not able to record all bijective relations, but instead can only record so called “collision free” relations. We prove that, with high probability, a bijection with random output is collision free, which allows us to perform path recording with slightly more error than incurred in [ma2024pseudorandom].
@@
$\mathsf{FHFHFO}\ket{x}\ket{\theta^{(1)},\theta^{(2)},\theta^{(3)}}=\mathsf{F}_{\theta^{(3)}}\cdot H^{\otimes n}\cdot\mathsf{F}_{\theta^{(2)}}\cdot H^{\otimes n}\cdot\mathsf{F}_{\theta^{(1)}}\ket{x}\ket{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\,.$
@@
$\frac{1}{2^{3n}}\sum_{\begin{subarray}{c}\theta_{1}^{(1)},\ldots,\theta_{2^{n}}^{(1)}\\ \theta_{1}^{(2)},\ldots,\theta_{2^{n}}^{(2)}\\ \theta_{1}^{(3)},\ldots,\theta_{2^{n}}^{(3)}\end{subarray}}\ket{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\,.$
@@
Since the view of the adversary does not contain the purifying register, when we trace them out we will get a uniform mixture of states corresponding to $\mathsf{FHFHF}$ for uniformly random angles $\theta^{(1)}$, $\theta^{(2)}$, and $\theta^{(3)}$. This corresponds exactly to choosing $\theta^{(1)}$, $\theta^{(2)}$, and $\theta^{(3)}$ uniformly at random. ∎
We now define a family of orthogonal states that correspond to a database $R$. We some new notation here to make the remaining equations less cumbersome. For a set $S$ of $n$-bit strings and $n$-bit string $z$, define $S_{z}$ to be:
@@
$\mathrm{Im}(R)$ =$=\{y:(x,y)\in R\}\,.$
As a first step, we define so-called $\mathsf{FHFHF}$-set states, as follows:
Definition 6.21 (FHFHF-set states).
Given a set $S$, define the set |$\ket{\mathsf{fhfhf}_{S}}$ as follows
$\frac{1}{2^{n}}\sum_{\theta_{1},\ldots,\theta_{2^{n}}\in[2^{n}]}\exp\left(i\sum_{z}\theta_{z}S_{z}\right)\ket{\theta}\,.$
The following lemma will show that these set states are orthogonal whenever the sets are different.
We show that these states form an orthogonal basis. In order to prove this, we first prove the following property, which comes from Fourier analysis of Boolean functions.
@@
Lemma 6.23 (Orthogonality of set states).
Set $S$, $S^{\prime}$ be two sets of $n$-bit strings, then the set states |$\ket{\mathsf{fhfhf}_{S}}$ and |$\ket{\mathsf{fhfhf}_{S^{\prime}}}$ are orthogonal.
Proof.
Let $z^{*}$ be the bit string on which $S_{z}$ and $S^{\prime}_{z}$ differ (given by Lemma 6.22). Then we have the following
$\braket{\mathsf{fhfhf}_{S}}{\mathsf{fhfhf}_{S^{\prime}}}$ $=\frac{1}{2^{2n}}\sum_{\theta_{1},\ldots,\theta_{2^{n}}}\left(i\sum_{z}\theta_{z}(S_{z}-S^{\prime}_{z})\right)$
$=\frac{1}{2^{2n}}\sum_{\theta_{z^{*}}}\exp\left(i\theta_{z^{*}}S_{z^{*}}-S^{\prime}_{z^{*}})\right)\cdot\sum_{\theta_{i}|i\neq z^{*}}\exp\left(i\sum_{z\neq z^{*}}\theta_{z}(S_{z}-S^{\prime}_{z})\right)$
$=0\,.$
Here we use the fact that the following equality:
$\mathbb{E}_{\theta}e^{i\theta x}=\begin{cases}1\quad x=0\\ 0\quad\text{otherwise}\end{cases}\,.$
Moreover, it is easy to check tht $\mathbb{E}_{k}e^{\mathrm{i}2\pi k/q}=\mathbb{E}_{\theta}e^{i\theta x}$ for an integer $x$ with $|x|<q$. Since $S_{z^{*}}-S^{\prime}_{z^{*}}$ is not 00, this term evalutes to 00 and therefore the entire product is 00. ∎
Definition 6.24 (FHFHF-relation states).
|$\ket{\mathsf{fhfhf}_{R}}=\frac{1}{2^{nt/2}}\sum_{w_{1},\ldots,w_{t}\in\{0,1\}^{n}}\left(\prod_{i}(-1)^{\braket{w_{i},y_{i}\oplus x_{i}}}\right)\\ \frac{1}{2^{3n}}\sum_{\begin{subarray}{c}\theta_{1}^{(1)},\ldots,\theta_{2^{n}}^{(1)}\\ \theta_{1}^{(2)},\ldots,\theta_{2^{n}}^{(2)}\\ \theta_{1}^{(3)},\ldots,\theta_{2^{n}}^{(3)}\end{subarray}}\exp\left(i\sum_{z,i}\theta^{(1)}_{z}(-1)^{\braket{z,x_{i}}}+\theta^{(2)}_{z}(-1)^{\braket{z,w_{i}}}+(-1)^{\braket{z,y_{i}}}\theta_{z}^{(3)}\right)\\ \ket{\theta^{(1)}}\otimes\ket{\theta^{(2)}}\otimes\ket{\theta^{(3)}}\,.$
Remark 6.25.
We note that in terms of the set states, we can re-write the state |$\ket{\mathsf{fhfhf}_{R}}$ as follows:
|$\ket{\mathsf{fhfhf}_{R}}=\ket{\mathsf{fhfhf}_{\mathrm{Dom}(R)}}\left(\frac{1}{2^{nt/2}}\sum_{w_{1},\ldots,w_{t}}\prod_{i}\left((-1)^{\braket{w_{i},x_{i}\oplus y_{i}}}\right)\ket{\mathsf{fhfhf}_{\{w_{1},\ldots,w_{t}\}}}\right)\ket{\mathsf{fhfhf}_{\mathrm{Im}(R)}}\,.$
Unlike in the analysis of the $\mathsf{PFC}$-ensemble, the $\mathsf{FHFHFC}$-relation states will not work on bijections, but on a slightly more restrictive set of relations, defined below.
Definition 6.26 (Collision-free relation).
Let $x_{1},\ldots x_{t}$ and $y_{1},\ldots,y_{t}$ be a collection of $n$-bit string. Then we say the pair is called “collision-free” if the following holds. For all permutations $\pi\in S_{t}$,
$\{x_{i}\oplus y_{i}\}\neq\{x_{i}\oplus y_{\pi(i)}\}\,.$
Given a set of $x_{1},\ldots,x_{t}$, let $\mathsf{CF}_{\mathbf{x}}$ be the set of collision free $y_{1},\ldots,y_{t}$.
Lemma 6.27 (Collision probability bound).
Fix distinct $x_{1},\ldots,x_{t}$, then the probability over uniformly chosen, distinct $y_{1},\ldots,y_{t}$ that the pair of sets is not collision free is at most $t^{3}/2^{n}$.
Proof.
We claim that being a collision-free relation implies that there exists a pair of $y_{i}$ and $y_{j}$ such that $y_{i}\oplus y_{j}$ is equal to $x_{1}\oplus x_{j}$ (here note that the second indices are the same). We will then upper bound the probability that any of the $t^{2}$ possible $y_{i}\oplus y_{j}$ (for randomly chosen $y_{i}$ and $y_{j}$) is equal to one of the at most $t$ many $x_{1}\oplus x_{j}$.
To see this, assume for the sake of contradiction that there is a permutation that breaks the collision-free property. Then it must be the case that $x_{1}\oplus y_{\pi(1)}=x_{a}\oplus y_{a}$ for some index $a$, we see that the following holds:
$x_{1}\oplus y_{1}\oplus x_{1}\oplus y_{\pi(1)}=x_{1}\oplus y_{1}\oplus x_{a}\oplus y_{a}$
$\implies y_{\pi(1)}\oplus y_{a}=x_{1}\oplus x_{a}\,.$
Since we assume that $x_{1},\ldots,x_{t}$ are fixed, the probability that the XOR of a single pair of $y_{i}$, $y_{j}$ fall in the set $\{x_{1}\oplus x_{j}:j\in[t]\}$ is at most $t/2^{n}$. By a union bound, the probability that any of them fall into this set is at most $t^{3}/2^{n}$. By this argument, if none of the XORs of $y_{i}$ and $y_{j}$ fall into the XOR of all pairs of $x$’s, the set must be collision free. ∎
Remark 6.28.
In general, one might hope that by not just considering collision free bijections, but any set of $x$’s and $y$’s for which the ordering can be determined uniquely by the list of XORs of the two tuples, one can remove the need for the Clifford unitary at the end. However, we leave this for future work.
@@
Let $R$ and $R^{\prime}$ be two different collision-free relations of size at most $2^{n-1}$ elements. Then |$\ket{\mathsf{fhfhf}_{R}}$ and |$\ket{\mathsf{fhfhf}_{R^{\prime}}}$ are orthogonal.
@@
From Lemma 6.23, together with Remark 6.25, if the image or domain of the two relations is different, then the states will be orthogonal since the set states storing the image and domain will differ. Therefore, we only have to consider pairs of collision free relations that have the same inputs and outputs. However, by assumption these are collision free. Say that these two relations are related by a permutation $\pi$ on the $y_{1},\ldots,y_{t}$, and say that $y_{i^{*}}$ is the index on which $x_{i}\oplus y_{i}\neq x_{i}\oplus y_{\pi(i)}$, which exists because of the collision free property.
$\braket{\mathsf{fhfhfh}_{R}}{\mathsf{fhfhfh}_{R^{\prime}}}$ $=\frac{1}{2^{nt}}\sum_{\begin{subarray}{c}w_{1},\ldots,w_{t}\\ w^{\prime}_{1},\ldots,w^{\prime}_{t}\end{subarray}}\prod_{i}(-1)^{\braket{w_{i},(x_{i}\oplus y_{i})}+\braket{w_{i}^{\prime},(x_{i}\oplus y_{\pi(i)})}}\braket{\mathsf{fhfhf}_{\{w_{1},\ldots,w_{t}\}}}{\mathsf{fhfhf}_{\{w^{\prime}_{1},\ldots,w^{\prime}_{t}\}}}$
$=\frac{1}{2^{nt}}\sum_{w_{1},\ldots,w_{t}}\prod_{i}(-1)^{\braket{w_{i},(x_{i}\oplus y_{i})+(x_{i}\oplus y_{\pi(i)})}}$
$=\frac{1}{2^{nt}}\left(\sum_{w_{i^{*}}}(-1)^{\braket{w_{i^{*}},(x_{i^{*}}\oplus(y_{i^{*}}+y_{\pi(i^{*})}))}}\right)\cdot\left(\sum_{\begin{subarray}{c}w_{1},\ldots,w_{i^{*}-1},\\ w_{i^{*}+1},\ldots,w_{t}\end{subarray}}\prod_{i\neq i^{*}}(-1)^{\braket{w_{i},(x_{i}\oplus y_{i})-(x_{i}\oplus y_{\pi(i)})}}\right)$
$=0\,.$
To get to the last line, we use the fact that in the first term, $x_{i^{*}}\oplus(y_{i^{*}}-y_{\pi_{i^{*}}})\neq 0$, so summing over the values of $w_{i^{*}}$, half will have a sign of $+1$ and half will have a sign of $-1$, yielding 00. ∎
@@
[…]
@@
|$\ket{x}\frac{1}{2^{3n}}\frac{1}{2^{3n}}\sum_{\begin{subarray}{c}\theta_{1}^{(1)},\ldots,\theta_{2^{n}}^{(1)}\\ \theta_{1}^{(2)},\ldots,\theta_{2^{n}}^{(2)}\\ \theta_{1}^{(3)},\ldots,\theta_{2^{n}}^{(3)}\end{subarray}}\alpha_{R}^{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}\right)\ket{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\,.$
@@
$\frac{1}{2^{n/2}}\sum_{y\in\{0,1\}^{n}}(-1)^{\braket{x,w}}\ket{w}\frac{1}{2^{3n}}\sum_{\begin{subarray}{c}\theta_{1}^{(1)},\ldots,\theta_{2^{n}}^{(1)}\\ \theta_{1}^{(2)},\ldots,\theta_{2^{n}}^{(2)}\\ \theta_{1}^{(3)},\ldots,\theta_{2^{n}}^{(3)}\end{subarray}}\alpha_{R}^{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}\right)\ket{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\,.$
@@
$\frac{1}{2^{n/2}}\sum_{w\in\{0,1\}^{n}}(-1)^{\braket{x,w}}\ket{w}\sum_{\begin{subarray}{c}\theta_{1}^{(1)},\ldots,\theta_{2^{n}}^{(1)}\\ \theta_{1}^{(2)},\ldots,\theta_{2^{n}}^{(2)}\\ \theta_{1}^{(3)},\ldots,\theta_{2^{n}}^{(3)}\end{subarray}}\alpha_{R}^{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\exp\left(i\sum_{z}\theta_{z}^{(1)}(-1)^{\braket{x,z}}+\theta_{z}^{(2)}(-1)^{\braket{w,z}}\right)\ket{\theta^{(1)},\theta^{(2)},\theta^{(3)}}\,.$